# Does rare, noise-induced, bypass transition in plane Couette flow bypass instantons ?

### v1

@@ 2.2.1 The algorithm
Figure 3: (a) Sketch of two bistable states $\mathcal{A}$ and $\mathcal{B}$ and the hypersurface $\mathcal{C}$ closely surrounding $\mathcal{A}$. two realisations of the dynamics are sketched: a single excursion in blue and a first passage trajectory in black and red. The red part of the first passage trajectory is the reactive trajectory (figure originally made for [Rolland and Simonnet, 2015]). (b) Sketch of the principle of AMS, showing to iterations of the algorithm, with $N=3$ clones, indicating the starting state $\mathcal{A}$ and its neighbourhood, the arrival state and its neighbourhood $\mathcal{B}$, three trajectories are ordered by their $\max_{t}\Phi$. Trajectory one (dashed blue line) is suppressed and branched on another trajectory at level $\max_{t}\Phi_{1}$ and then ran according to its natural dynamics. Trajectory $2$ is then suppressed and branched on 3 at level $\max_{t}\Phi_{2}$ (figure originally made for [Simonnet, 2016]).
@@
In order to present the numerical method used to compute rare trajectories, let us first give a formal phase space description of the rare events we will study in this text. We sketch the build up trajectories in figure 3 (a) and consider $\mathcal{A}$ a neighbourhood of the laminar flow in phase space, an hypersurface $\mathcal{C}$ that closely surrounds $\mathcal{A}$ and $\mathcal{B}$ a neighbourhood of the turbulent flow in phase space. A realisation of the dynamics which starts in $\mathcal{A}$ fluctuates around it (for instance has several excursions out of $\mathcal{C}$, an hypersurface closely surrounding $\mathcal{A}$) and eventually crosses $\mathcal{C}$ and reaches $\mathcal{B}$ before coming back to $\mathcal{A}$ is termed a first passage. Its average duration is termed the mean first passage time $T$. The last stage of the dynamics is termed a reactive trajectory: this is the part of the dynamics that starts in $\mathcal{A}$, crosses $\mathcal{C}$ and reaches $\mathcal{B}$ before $\mathcal{A}$. Quantitative definitions of sets $\mathcal{A}$, $\mathcal{B}$ and hypersurface $\mathcal{C}$ for our case will be given in section 2.2.2.
@@
We then present the generic formulation for two of the variants of Adaptive Multilevel Splitting (AMS) which were used for the computation presented in this text They were initially proposed by Cérou & Guyader [Cérou and Guyader, 2007]. Both variants of AMS use a reaction coordinate (or observable) $\phi(\mathbf{u}(t))=\Phi(t)$. The reaction coordinate $\phi$ is a mapping from $\mathbb{R}^{D}$, where the discretised velocity field $\mathbf{u}$ lives, to $[0,1]$. The reaction coordinate can be viewed as a function of time $\Phi(t)$ on trajectories. The reaction coordinate measures the position of the flow relatively to starting set $\mathcal{A}$, where is it 00 and arrival set $\mathcal{B}$ where it is $1$. We will define our reaction coordinate quantitatively in section 2.2.2. All variants of AMS run $N$ clone dynamics of the system to compute iteratively from $N-N_{c}+1$ up to $N$ reactive trajectories going from set $\mathcal{A}$ to the hypersurface $\mathcal{C}$, closely surrounding set $\mathcal{A}$, then from $\mathcal{C}$ to the set $\mathcal{B}$. Here $1\leq N_{c}<N$ is the number of trajectories that will be removed at each iteration. We will label the velocity field and pressure field of each trajectory by $1\leq i\leq N$ as in $\left(\begin{matrix}\mathbf{u}_{i}\\ p_{i}\end{matrix}\right)(t)$. The reaction coordinate as a function of time for the corresponding trajectory is then denoted by $\Phi_{i}(t)=\phi(\mathbf{u}_{i}(t))$. The algorithm is sketched in figure 3 (b) and proceeds in the following manner
@@
  • The algorithm then runs the stages of mutation selection. At each stage, the clones $1\leq i\leq N$, are ordered by $\max_{t}\Phi_{i}(t)$: the maximum of the reaction coordinate achieved on the trajectory. The $N_{c}$ clones $1\leq i^{\prime}\leq N_{c}$ having the smallest values of $\max_{t}\Phi_{i^{\prime}}(t)$ are removed. In order to keep a constant number of clones, $N_{c}$ new clones are generated by branching. We draw $N_{c}$ clones $i^{\prime\prime}$ uniformly from the clones $N_{c}+1\leq i\leq N$ such that $\max_{t}\Phi_{i}(t)\geq\max_{t}\Phi_{N_{c}}(t)$. We denote by $t_{i^{\prime\prime}}$ the first time such that clone $i^{\prime\prime}$ reaches the reaction coordinate value $\max_{t}\Phi_{i^{\prime}}(t)$: $t_{i^{\prime\prime}}=\min\left\{t^{\prime}\geq 0\backslash\Phi_{i^{\prime\prime}}(t^{\prime})>\max_{t}\Phi_{i^{\prime}}(t)\right\}$. The new clone $i^{\prime}$ is such that $\mathbf{u}_{i^{\prime}}(t)=\mathbf{u}_{i^{\prime\prime}}(t)$, $p_{i^{\prime}}(t)=p_{i^{\prime\prime}}(t)$, for $0\leq t\leq t_{i^{\prime\prime}}$. For $t>t_{i^{\prime\prime}}$ the clone $i^{\prime}$ then follows its natural dynamics under a new, independent realisation of the forcing $\mathbf{f}$ until it reaches either $\mathcal{A}$ or $\mathcal{B}$.
@@
The algorithm stops after iteration number $\kappa\in\mathbb{N}$, the first iteration where at least $N-N_{c}+1$ clones trajectories have reached $\mathcal{B}$. We denote by $r\in\left[1-\frac{N_{c}-1}{N},1\right]$ the proportion of clones that have reached $\mathcal{B}$ at the last stage. Note that $\kappa$ and $r$ are a priori random numbers, different from one AMS run to the other, whose value depend on the independent noise realisations drawn during simulation of trajectories. This yields an estimator of the probability $\alpha$ of reaching $\mathcal{B}$ before $\mathcal{A}$ starting from the distribution of first exit points on $\mathcal{C}$, and the corresponding mean first passage time $T$
@@
$\alpha=r\left(1-\frac{N_{c}}{N}\right)^{\kappa}\,,\,\hat{\alpha}=\left\langle\alpha\right\rangle_{o}\,,\,\hat{T}=\left\langle\left(\frac{1}{\alpha}-1\right)\tilde{\tau}+\tau\right\rangle_{o}\,,$ (10)
@@
where $\tau$ is the mean duration of reactive trajectories and $\tilde{\tau}$ is the mean duration of non reactive trajectories, computed in each AMS run and $\langle\cdot\rangle_{o}$ corresponds to an arithmetic average over $o$ AMS runs [Cérou and Guyader, 2007, Rolland, 2018]. In this text, we both performed calculations with $N_{c}=1$ (systematically leading to $r=1$) and $N_{c}>1$. The algorithm is naturally parallelised over the $N_{c}$ suppressed clones: the speed up using $N_{p}$ cores has been measured (not shown here) and requires $N_{c}/N_{p}$ to be strictly larger than one to remain close to linear.
@@
As it was done in other studies [Simonnet et al., 2021, Rolland, 2022], we save a state at the last step $\kappa$ of the AMS computation. This is the state where the trajectory realising the largest $\Phi$ among the last non passing reactive trajectory suppressed is branched. We use the same terminology as in [Rolland, 2022], we term it the last stage at the last step and denote it by $\mathbf{u}_{\rm last}$. Using the notations introduced above, this state is computed using the clone $\mathbf{u}_{i^{\prime\prime}}$ that replaces $\mathbf{u}_{i^{\prime}=N_{c}}$. We then have $\mathbf{u}_{\rm last}=\mathbf{u}_{i^{\prime\prime}}(t_{i^{\prime\prime}}$). In complex spatially extended systems, it has been verified that this state is very similar to the saddle crossed by the trajectory [Simonnet et al., 2021]. In gradient systems, (with few DoF or with spatial extension [Rolland et al., 2016]), it can be verified that this gives a finer and finer approximation of the actual, perfectly defined saddle as the noise variance is decreased (not shown here). In unforced system, this indicates a turning point for the trajectory [Rolland, 2022].
@@ 2.2.2 Reaction coordinate
Because the kinetic energy $E_{k}$ is almost 00 for the laminar state, takes its largest possible values when the flow is turbulent and grows in between, we used a reaction coordinate affine in spatially averaged kinetic energy
@@
This reaction coordinate is parameterised by two specific values of kinetic energy $E_{\mathcal{A}}$ and $E_{\mathcal{B}}$ which depend indirectly on the control parameters used in the AMS calculations: the Reynolds number, the energy injection rate and the shape of the noise spectrum. The value $E_{\mathcal{A}}$ should be such that $E_{k}(t)\simeq E_{\mathcal{A}}$ and $\Phi(t)\simeq 0$ when the flow fluctuates near the laminar state under forcing. The exact value of $E_{\mathcal{A}}$ is computed using time series of kinetic energy sampled at the control parameters of interest: $E_{\mathcal{A}}$ is defined as the average of $E_{k}(t)$ conditioned on no transition occurring. The other parameter $E_{\mathcal{B}}$ is such that the reaction coordinate is one when the flow has reached the neighbourhood of the turbulent state. Equation (11) then implies that $E_{\mathcal{B}}+E_{\mathcal{A}}\simeq E_{k,\text{turb}}$. Since $E_{\mathcal{A}}$ will often be relatively small, we thus chose $E_{\mathcal{B}}=0.06$ regardless of control parameters, aside of a few specified cases where we used $E_{\mathcal{B}}=0.06-E_{\mathcal{A}}$.
@@
where $\sigma$ is the standard deviation of $E_{k}(t)$ conditioned on no transition occurring, sampled from the same dataset as $E_{\mathcal{A}}$. The rationale is that the flow leaves the neighbourhood of the laminar state when $E_{k}\geq E_{\mathcal{A}}+\sigma$.
@@
While some AMS computations used ad hoc parameters sampled on a case by case basis, for most values of the control parameters $R$ and $\epsilon$, We did not resample $E_{k}(t)$ to compute $E_{\mathcal{A}}$ and $\sigma$. Instead, we used the parametric dependence of $E_{\mathcal{A}}$ and $\sigma$ on $\beta$ to interpolate values. Indeed, we can note that provided $\beta$ is large enough, or conversely $\epsilon$ small enough (the two being related, as seen in (Eq. (8))) for build up events to be rare, we find that $E_{\mathcal{A}}$ and $\sigma$ are proportional to $\frac{1}{\beta}$. Moreover, for the two distinct families of forcing, we find that $E_{\mathcal{A}}$ and $\sigma$ are proportional to $\epsilon$ in good approximation (Fig. 4 (a) for forcing on component and Fig. 4 (b) for divergence free forcing). This parametric dependence is in line with upper bounds of kinetic energy found for the linearised flow [Farrell and Ioannou, 1993b]. Note however that the coefficient of this linear dependence depend on the main class of forcing. The proportionality coefficient in the divergence free case is larger than the one is the case of forcing on components by a factor more than 10. We also note that within a class of forcing we do not get exactly the same proportionality coefficient between $E_{\mathcal{A}}$ and $\epsilon$. While $E_{\mathcal{A}}$ and $\sigma$ grow with the energy injection rate, this simplified dependence is an approximation. Owing to the different spectrum shape and the variety of process bringing energy from the forcing to the receptive scale (direct forcing and transfer across scales), the actual dependence is actually more complex, and the relation between $E_{\mathcal{A}}$ and $\epsilon$ is not universal. We finally note that changing the numerical resolution does not lead to a change in the way $E_{\mathcal{A}}$ and $\sigma$ depend on $\epsilon$. Note that in more resolved simulations, we have pushed the cut-off of the spectrum to larger scales, so that the forcing spectrum is different. Because of that and because the forcing is white in the wall normal direction, a higher resolved simulation receives more energy for the same value of $\beta$, as expected from equation (8). We have also tested this relation with a larger cut-off size of $L_{c}=6$ (not shown here), which leads to a forcing that can directly affect the most energetic scales. The same linear dependence of $E_{\mathcal{A}}$ on $\frac{1}{\epsilon}$ is found, but in that case the values of $E$ are slightly above those computed when $L_{c}=1$ in the forcing, possibly because large scales now do not only receive energy from non linear redistribution.
@@
For this matter, for $M$ values of $\beta_{a}\in\left[4\cdot 10^{5},10^{10}\right]$, $1\leq a\leq M$, we sample $E_{a}$ and $\sigma_{a}$ in time series where the flow does not transit to turbulence (this is extremely rare except maybe for $\beta=4\cdot 10^{5}$). Linear regression gives
$d_{E}=\frac{M\sum_{a=1}^{M}(\log(E_{a})\log(\beta_{a}))-\left(\sum_{a=1}^{M}\log(E_{a})\right)\left(\sum_{a=1}^{M}\log(\beta_{a})\right)}{M\sum_{a=1}^{M}(\log(\beta_{a}))^{2}-\left(\sum_{a=1}^{M}\log(\beta_{a})\right)^{2}}\,,$ (14)
$o_{E}=\frac{\left(\sum_{a=1}^{M}\log(\beta_{a})\right)\left(\sum_{a=1}^{M}\log(\beta_{a})\log(E_{a})\right)-\left(\sum_{a=1}^{M}\log(E_{a})\right)\left(\sum_{a=1^{M}}\log(\beta_{a})^{2}\right)}{\left(\sum_{a=1}^{M}\log(\beta_{a})\right)^{2}-M\sum_{a=1}^{M}\log(\beta_{a})^{2}}\,,$ (15)
$d_{S}=\frac{M\sum_{a=1}^{M}(\log(\sigma_{a})\log(\beta_{a}))-\left(\sum_{a=1}^{M}\log(\sigma_{a})\right)\left(\sum_{a=1}^{M}\log(\beta_{a})\right)}{M\sum_{a=1}^{M}(\log(\beta_{a}))^{2}-\left(\sum_{a=1}^{M}\log(\beta_{a})\right)^{2}}\,,$ (16)
$o_{S}=\frac{\left(\sum_{a=1}^{M}\log(\beta_{a})\right)\left(\sum_{a=1}^{M}\log(\beta_{a})\log(\sigma_{a})\right)-\left(\sum_{a=1}^{M}\log(\sigma_{a})\right)\left(\sum_{a=1}^{M}\log(\beta_{a})^{2}\right)}{\left(\sum_{a=1}^{M}\log(\beta_{a})\right)^{2}-M\sum_{a=1}^{M}\log(\beta_{a})^{2}}\,.$ (17)
We always have $d_{E}\simeq d_{S}\simeq-1$, owing to the proportionality relation between $E_{\mathcal{A}}$, $\sigma$ and $\frac{1}{\beta}$ (as reported in table 1, § B). Using equation (13), we can interpolate/extrapolate $\log(E_{\mathcal{A}})$ and $\log(\sigma)$ for most values of $\beta$, we then exponentiate.
We give the value used to define the sets $\mathcal{A}$, $\mathcal{C}$ and $\mathcal{B}$ for each set of simulations in appendix B. This comprises of specific cases where we will give $E_{\mathcal{A}}$, $E_{\mathcal{B}}$ and $\sigma$ and generic cases where we give $d_{E}$, $o_{E}$, $d_{S}$ and $o_{S}$, which mostly concern simulations performed in the largest systems.

### v2

@@ 2.2.1 The algorithm
Figure 3: (a) Sketch of two bistable states $\mathcal{A}$ and $\mathcal{B}$ and the hypersurface $\mathcal{C}$ closely surrounding $\mathcal{A}$. two realisations of the dynamics are sketched: a single excursion in blue and a first passage trajectory in black and red. The red part of the first passage trajectory is the reactive trajectory (figure originally made for Rolland and Simonnet, (2015)). (b) Sketch of the principle of AMS, showing to iterations of the algorithm, with $N=3$ clones, indicating the starting state $\mathcal{A}$ and its neighbourhood, the arrival state and its neighbourhood $\mathcal{B}$, three trajectories are ordered by their $\max_{t}\Phi$. Trajectory one (dashed blue line) is suppressed and branched on another trajectory at level $\max_{t}\Phi_{1}$ and then ran according to its natural dynamics. Trajectory $2$ is then suppressed and branched on 3 at level $\max_{t}\Phi_{2}$ (figure originally made for Simonnet, (2016)).
@@
In order to present the numerical method used to compute rare trajectories, let us first give a formal phase space description of the rare events we will study in this text. We sketch the build up trajectories in figure 3 (a) and consider $\mathcal{A}$ a neighbourhood of the laminar flow in phase space, an hypersurface $\mathcal{C}$ that closely surrounds $\mathcal{A}$ and $\mathcal{B}$ a neighbourhood of the turbulent flow in phase space. A realisation of the dynamics which starts in $\mathcal{A}$ fluctuates around it (for instance has several excursions out of $\mathcal{C}$, an hypersurface closely surrounding $\mathcal{A}$) and eventually crosses $\mathcal{C}$ and reaches $\mathcal{B}$ before coming back to $\mathcal{A}$ is termed a first passage (Black then red curve in Fig. 3 (a)). Its average duration is termed the mean first passage time $T$. The last part of the first passage is termed a reactive trajectory: this is the part of the dynamics that starts in $\mathcal{A}$, crosses $\mathcal{C}$ and reaches $\mathcal{B}$ before $\mathcal{A}$ (Portion of black curve between $\mathcal{A}$ and $\mathcal{C}$ continued by red curve in Fig. 3 (a)). Quantitative definitions of sets $\mathcal{A}$, $\mathcal{B}$ and hypersurface $\mathcal{C}$ for our case will be given in section 2.2.2.
@@
We then present the generic formulation for two of the variants of Adaptive Multilevel Splitting (AMS) which were used for the computation presented in this text They were initially proposed by Cérou & Guyader Cérou and Guyader, (2007). Both variants of AMS use a reaction coordinate (or observable) $\phi(\mathbf{u}(t))=\Phi(t)$. The reaction coordinate $\phi$ is a mapping from $\mathbb{R}^{D}$, where the discretised velocity field $\mathbf{u}$ lives, to $[0,1]$. The reaction coordinate can be viewed as a function of time $\Phi(t)$ on trajectories. The reaction coordinate measures the position of the flow relatively to starting set $\mathcal{A}$, where is it 00 and arrival set $\mathcal{B}$ where it is $1$. We will define our reaction coordinate quantitatively in section 2.2.2. All variants of AMS run $N$ clone dynamics of the system to compute iteratively from $N-N_{c}+1$ up to $N$ reactive trajectories going from set $\mathcal{A}$ to the hypersurface $\mathcal{C}$, closely surrounding set $\mathcal{A}$, then from $\mathcal{C}$ to the set $\mathcal{B}$. Here $1\leq N_{c}<N$ is the number of trajectories that will be removed at each iteration. We will label the velocity field and pressure field of each trajectory by $1\leq i\leq N$ as in $\left(\begin{matrix}\mathbf{u}_{i}\\ p_{i}\end{matrix}\right)(t)$. The reaction coordinate as a function of time for the corresponding trajectory is then denoted by $\Phi_{i}(t)=\phi(\mathbf{u}_{i}(t))$. The algorithm is sketched in figure 3 (b) and proceeds in the following manner
@@
  • The algorithm then runs the steps of mutation selection. At each step, the clones $1\leq i\leq N$, are ordered by $\max_{t}\Phi_{i}(t)$: the maximum of the reaction coordinate achieved on the trajectory. The $N_{c}$ clones $1\leq i^{\prime}\leq N_{c}$ having the smallest values of $\max_{t}\Phi_{i^{\prime}}(t)$ are removed. In order to keep a constant number of clones, $N_{c}$ new clones are generated by branching. We draw $N_{c}$ clones $i^{\prime\prime}$ uniformly from the clones $N_{c}+1\leq i\leq N$ such that $\max_{t}\Phi_{i}(t)\geq\max_{t}\Phi_{N_{c}}(t)$. We denote by $t_{i^{\prime\prime}}$ the first time such that clone $i^{\prime\prime}$ reaches the reaction coordinate value $\max_{t}\Phi_{i^{\prime}}(t)$: $t_{i^{\prime\prime}}=\min\left\{t^{\prime}\geq 0\backslash\Phi_{i^{\prime\prime}}(t^{\prime})>\max_{t}\Phi_{i^{\prime}}(t)\right\}$. The new clone $i^{\prime}$ is such that $\mathbf{u}_{i^{\prime}}(t)=\mathbf{u}_{i^{\prime\prime}}(t)$, $p_{i^{\prime}}(t)=p_{i^{\prime\prime}}(t)$, for $0\leq t\leq t_{i^{\prime\prime}}$. For $t>t_{i^{\prime\prime}}$ the clone $i^{\prime}$ then follows its natural dynamics under a new, independent realisation of the forcing $\mathbf{f}$ until it reaches either $\mathcal{A}$ or $\mathcal{B}$.
@@
The algorithm stops after iteration number $\kappa\in\mathbb{N}$, the first iteration where at least $N-N_{c}+1$ clones trajectories have reached $\mathcal{B}$. We denote by $r\in\left[1-\frac{N_{c}-1}{N},1\right]$ the proportion of clones that have reached $\mathcal{B}$ at the last step. Note that $\kappa$ and $r$ are a priori random numbers, different from one AMS run to the other, whose value depend on the independent noise realisations drawn during simulation of trajectories. This yields an estimator of the probability $\alpha$ of reaching $\mathcal{B}$ before $\mathcal{A}$ starting from the distribution of first exit points on $\mathcal{C}$, and the corresponding mean first passage time $T$
@@
$\alpha=r\left(1-\frac{N_{c}}{N}\right)^{\kappa}\,,\,\hat{\alpha}=\left\langle\alpha\right\rangle_{o}\,,\,\hat{T}=\left\langle\left(\frac{1}{\alpha}-1\right)\tilde{\tau}+\tau\right\rangle_{o}\,,$
@@
where $\tau$ is the mean duration of reactive trajectories and $\tilde{\tau}$ is the mean duration of non reactive trajectories, computed in each AMS run and $\langle\cdot\rangle_{o}$ corresponds to an arithmetic average over $o$ AMS runs Cérou and Guyader, (2007); Rolland, (2018). In particular, note that the expression of the estimator of $\alpha$, is a consequence of the fact that the estimate probability to transit from one branching level to the next is $1-\frac{N_{c}}{N}$. Moreover, the estimated probability to go from the last branching level to $\mathcal{B}$ is one. In this text, we both performed calculations with $N_{c}=1$ (systematically leading to $r=1$) and $N_{c}>1$. The algorithm is naturally parallelised over the $N_{c}$ suppressed clones: the speed up using $N_{p}$ cores has been measured (not shown here) and requires $N_{c}/N_{p}$ to be strictly larger than one to remain close to linear. Note that other procedures used to compute instantons use intermediate states Lecoanet and Kerswell, (2018). In that case, all the intermediate states are found on the fluctuation part of the trajectory. This is similar to what is found in AMS computations, since the probability to go from the saddle to state $\mathcal{B}$ goes to one as quantities controlling the rarity of the event like $\epsilon$ go to zero.
@@
As it was done in other studies Simonnet et al., (2021); Rolland, (2022), we save a state at the last step $\kappa$ of the AMS computation. This state is taken on the trajectory that has the largest reaction coordinate $\Phi$ among the remaining non reactive trajectories at the last step of an AMS computation. On this trajectory, this state is taken at branching time. We use the same terminology as in Rolland, (2022), we term it the last state at the last step and denote it by $\mathbf{u}_{\rm last}$. Using the notations introduced above, this state is computed using the clone $\mathbf{u}_{i^{\prime\prime}}$ that replaces $\mathbf{u}_{i^{\prime}=N_{c}}$. We then have $\mathbf{u}_{\rm last}=\mathbf{u}_{i^{\prime\prime}}(t_{i^{\prime\prime}}$). In complex spatially extended systems, it has been verified that this state is very similar to the saddle crossed by the trajectory Simonnet et al., (2021). In gradient systems, (with few DoF or with spatial extension Rolland et al., (2016)), it can be verified that this gives a finer and finer approximation of the actual, perfectly defined saddle as the noise variance is decreased (not shown here). In unforced system, this indicates a turning point for the trajectory Rolland, (2022). This is directly connected to the fact that the probability to transit from the last branching level to $\mathcal{B}$ is closer and closer to one as $\epsilon\rightarrow 0$. This means that the last branching level is then very close to the first instant of the relaxation path, which is said saddle in forced systems, in the limit $\epsilon\rightarrow 0$.
@@ 2.2.2 Reaction coordinate
We will construct a reaction coordinate which is monotonous in the kinetic energy $E_{k}$. Indeed, $E_{k}$ is almost 00 when the forced flow fluctuates around the first metastable state, the laminar baseflow. Meanwhile, it takes its largest possible values when the flow is turbulent (the second metastable state). Finally, $E_{k}$ grows when the forced flow travels from the laminar state to turbulence. For simplicity, we used a reaction coordinate affine in spatially averaged kinetic energy
@@
This reaction coordinate is parameterised by two specific values of kinetic energy $E_{\mathcal{A}}$ and $E_{\mathcal{B}}$ which depend indirectly on the control parameters used in the AMS calculations: the Reynolds number, the energy injection rate and the shape of the noise spectrum. The value $E_{\mathcal{A}}$ should be such that $E_{k}(t)\simeq E_{\mathcal{A}}$ and $\Phi(t)\simeq 0$ when the flow is near the first metastable state and fluctuates near the laminar state under the forcing. The exact value of $E_{\mathcal{A}}$ is computed using time series of kinetic energy sampled at the control parameters of interest: $E_{\mathcal{A}}$ is defined as the average of $E_{k}(t)$ conditioned on no transition occurring. The other parameter $E_{\mathcal{B}}$ is such that the reaction coordinate is one when the flow has reached the neighbourhood of the turbulent state. Equation (11) then implies that $E_{\mathcal{B}}+E_{\mathcal{A}}\simeq E_{k,\text{turb}}$. The conditional average of the kinetic energy $E_{\mathcal{A}}$ will often be relatively small: the energy injection rate is small enough for our description of the forced flow as two metastable states makes sense. As a consequence, we chose $E_{\mathcal{B}}=0.06$ regardless of control parameters, aside of a few specified cases where we used $E_{\mathcal{B}}=0.06-E_{\mathcal{A}}$.
@@
where $\sigma$ is the standard deviation of $E_{k}(t)$ conditioned on no transition occurring, sampled from the same dataset as $E_{\mathcal{A}}$. The rationale is that we have a rare occurrence of a reactive trajectory when the flow leaves the neighbourhood of the laminar state, that is to say when when $E_{k}\geq E_{\mathcal{A}}+\sigma$. The narrow region of phase space such that $E_{k}\leq E_{\mathcal{A}}+\sigma$ can then be viewed as the neighbourhood of the first metastable state, as long as $\sigma$ is small enough.
@@
While some AMS computations used ad hoc parameters sampled on a case by case basis, for most values of the control parameters $R$ and $\epsilon$, We did not resample $E_{k}(t)$ to compute $E_{\mathcal{A}}$ and $\sigma$. Instead, we used the parametric dependence of $E_{\mathcal{A}}$ and $\sigma$ on $\beta$ to interpolate values. Indeed, we can note that provided $\beta$ is large enough, or conversely $\epsilon$ small enough (the two being related, as seen in (Eq. (8))) for build up events to be rare, we find that $E_{\mathcal{A}}$ and $\sigma$ are proportional to $\frac{1}{\beta}$. Moreover, for the two distinct families of forcing, we find that $E_{\mathcal{A}}$ and $\sigma$ are proportional to $\epsilon$ in good approximation (Fig. 4 (a) for forcing on component and Fig. 4 (b) for divergence free forcing). This parametric dependence is in line with upper bounds of kinetic energy found for the linearised flow Farrell and Ioannou, 1993b . Note however that the coefficient of this linear dependence depend on the main class of forcing. The proportionality coefficient in the divergence free case is larger than the one is the case of forcing on components by a factor more than 10. We also note that within a class of forcing we do not get exactly the same proportionality coefficient between $E_{\mathcal{A}}$ and $\epsilon$. While $E_{\mathcal{A}}$ and $\sigma$ grow with the energy injection rate, this simplified dependence is an approximation. Owing to the different spectrum shape and the variety of process bringing energy from the forcing to the receptive scale (direct forcing and transfer across scales), the actual dependence is actually more complex, and the relation between $E_{\mathcal{A}}$ and $\epsilon$ is not universal. We finally note that changing the numerical resolution does not lead to a change in the way $E_{\mathcal{A}}$ and $\sigma$ depend on $\epsilon$. Note that in more resolved simulations, we have pushed the cut-off of the spectrum to larger scales, so that the forcing spectrum is different. Because of that and because the forcing is white in the wall normal direction, a higher resolved simulation receives more energy for the same value of $\beta$, as expected from equation (8). We have also tested this relation with a larger cut-off size of $L_{c}=6$ (not shown here), which leads to a forcing that can directly affect the most energetic scales. The same linear dependence of $E_{\mathcal{A}}$ on $\frac{1}{\epsilon}$ is found, but in that case the values of $E$ are slightly above those computed when $L_{c}=1$ in the forcing, possibly because large scales now do not only receive energy from non linear redistribution.
@@
For this matter, for $M$ values of $\beta_{a}\in\left[4\cdot 10^{5},10^{10}\right]$, $1\leq a\leq M$, we sample $E_{a}$ and $\sigma_{a}$ in time series where the flow does not transit to turbulence (this is extremely rare except maybe for $\beta=4\cdot 10^{5}$). The factors $d_{E}$, $o_{E}$, $d_{S}$ and $o_{S}$ are obtained using the explicit formula for linear regression by affine functions (Eq. (13)) by least squares using the $M$ sampled pairs of values (Fig. 4). We always have $d_{E}\simeq d_{S}\simeq-1$, owing to the proportionality relation between $E_{\mathcal{A}}$, $\sigma$ and $\frac{1}{\beta}$ (as reported in table 4, § C). Using equation (13), we can interpolate/extrapolate $\log(E_{\mathcal{A}})$ and $\log(\sigma)$ for most values of $\beta$, we then exponentiate.
We give the value used to define the sets $\mathcal{A}$, $\mathcal{C}$ and $\mathcal{B}$ for each set of simulations in appendix C. This comprises of specific cases where we will give $E_{\mathcal{A}}$, $E_{\mathcal{B}}$ and $\sigma$ and generic cases where we give $d_{E}$, $o_{E}$, $d_{S}$ and $o_{S}$, which mostly concern simulations performed in the largest systems. We finally note that in all considered case, the values of $E_{\mathcal{A}}$ and $\sigma$ are small enough for the evolution of the forced flow near the laminar base flow as small fluctuations around one metastable state.